# Configuration spaces of circles in the plane

### v1

@@ 5. Fundamental groups of configuration spaces of circles
In this section we prove our main results. First we focus on the much more difficult unlabeled case.
@@ 5.1. The unlabeled case
Let $T$ be a tree with $n$ non-root vertices. Then $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ is aspherical, and its fundamental group is isomorphic to $\operatorname{BAut}(T)$, hence it is a $K(\operatorname{BAut}(T),1)$.
@@
We should mention that sometimes the definition of “$K(G,1)$” includes a requirement that the space be homotopy equivalent to a CW-complex. It is clear that our spaces have this property. Indeed, $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$ is semialgebraic by Proposition 2.4 and hence triangulable by [Hir75], and $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ is the orbit space of this under a linear action of the finite group $S_{n}$, hence is triangulable, for example by [PS85]. The same therefore holds for each connected component $\operatorname{LConf}_{T}(S^{1},\mathbb{R}^{2})$ and $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$.
@@
Let $D$ be an open disk in the plane. Then the inclusion $D\to\mathbb{R}^{2}$ induces a homotopy equivalence from $\operatorname{Conf}_{n}(S^{1},D)$ to $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$. For $T$ a tree with $n$ non-root vertices, setting $\operatorname{Conf}_{T}(S^{1},D)\coloneqq\operatorname{Conf}_{n}(S^{1},D)\cap\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$, we also have a homotopy equivalence from $\operatorname{Conf}_{T}(S^{1},D)$ to $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$. Analogous statements hold in the labeled case.
@@
First, given $S,S^{\prime}\subseteq\mathbb{R}^{2}$, say that $S$ is barely contained in $S^{\prime}$ if $S$ is contained in $S^{\prime}$ but no $\varepsilon$-neighborhood of $S$ is contained in $S^{\prime}$ for any $\varepsilon>0$. If $S$ is a bounded subset of the plane, then there exists a unique closed disk $D^{\prime}$ in the plane such that $S$ is barely contained in $D^{\prime}$. Hence, for any bounded $S$ and any disk $D^{\prime}$, there is a unique way to scale and translate $S$ until the unique disk it is barely contained in is $D^{\prime}$.
Now let $\Phi\colon\operatorname{Conf}_{n}(S^{1},D)\to\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ be the map that sends $\{C_{1},\dots,C_{n}\}$ to itself, induced by the inclusion $D\to\mathbb{R}^{2}$. Let $\Psi\colon\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})\to\operatorname{Conf}_{n}(S^{1},D)$ be the map that takes a configuration $\{C_{1},\dots,C_{n}\}$ in the plane, and scales and translates it until $C_{1}\cup\cdots\cup C_{n}$ is barely contained in the closed disk of radius $r/2$ centered at $\vec{x}$, where $r$ and $\vec{x}$ are the radius and center of the open disk $D$. It is clear that $\Psi\circ\Phi$ and $\Phi\circ\Psi$ are homotopic to the relevant identity maps, since we can interpolate the scaling and translating continuously. The same argument works in the presence of a tree $T$ and/or in the labeled case, since scaling and translating does not change the property of one circle being nested in another, nor any chosen labeling of the circles. ∎
A key tool in the proof of Theorem A will be fiber bundles. Background can be found for example in [Hat02, Pages 375–377].
Definition 5.3 (Fiber bundle).
Let $\phi\colon E\to B$ be a continuous surjection of topological spaces and $F$ a topological space. We call $E$ a fiber bundle over $B$ (with respect to $\phi$, and with fiber $F$) if for all $b\in B$ there exists an open neighborhood $U$ of $b$ and a homeomorphism $h\colon\phi^{-1}(U)\to U\times F$ such that the projection $U\times F\to U$ composed with $h$ equals the restriction of $\phi$ to $\phi^{-1}(U)$. Equivalently, there exists a map $\psi\colon\phi^{-1}(U)\to F$ such that $\phi\times\psi$ restricts to a homeomorphism $\phi^{-1}(U)\to U\times F$. Intuitively, $E$ being a fiber bundle over $B$ with fiber $F$ means that $E$ locally resembles $B\times F$.
Note that for each $b\in B$ (and any choice of $U$), the homeomorphism $h$ must take $\phi^{-1}(b)$ homeomorphically to $F_{b}\coloneqq\{b\}\times F$. In particular, all the $\phi^{-1}(b)$ are homeomorphic to each other, and to $F$. If we have a fixed basepoint $b_{0}\in B$, we will often identify $F$ with $F_{b_{0}}$, and so refer to $F$ as a subspace of $E$.
@@
Now we can prove our main result, that $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ is aspherical and its fundamental group is isomorphic to $\operatorname{BAut}(T)$.
@@
If $n=0$ then the theorem follows from the trivial fact that a single point is contractible. Now suppose $n\geq 1$.
Construction of the fiber bundle: Given an (unlabeled) configuration $\{C_{1},\dots,C_{n}\}$ in $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$, call $C_{i}$ a top circle if it is not nested in any $C_{j}$, i.e, if $\mathbb{R}^{2}\vdash C_{i}$. The number of top circles in any configuration in $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ equals the number of children of the root $\varnothing$ of $T$, call this $m$. Since $n\geq 1$, we know $m\geq 1$. Let $\Lambda$ be the induced subtree of $T$ consisting of just the root and its children. If $\{v_{1},\dots,v_{m}\}$ is the set of children of the root, then $T$ is the union of $\Lambda$ with all the $T(v_{i})$ for $1\leq i\leq m$.
Define a surjective map
$\phi\colon\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})\to\operatorname{Conf}_{\Lambda}(S^{1},\mathbb{R}^{2})$
by sending a configuration to the configuration consisting only of its top circles. Set $b_{0}=\varkappa_{\Lambda}$ and $x_{0}=\varkappa_{T}$, so $\phi(x_{0})=b_{0}$, and let $F$ consist of all configurations in $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ whose top circles are precisely $\mathfrak{C}_{1}^{\Lambda},\dots,\mathfrak{C}_{m}^{\Lambda}$, so $F$ is the fiber of $b_{0}$. Identify $F$ with its image in $\operatorname{Conf}_{n-m}(S^{1},\mathbb{R}^{2})$ obtained by deleting these top circles. Writing $\mathfrak{D}_{i}^{\Lambda}$ for the open disk whose boundary is $\mathfrak{C}_{i}^{\Lambda}$, under this identification the connected component of $\varkappa_{T}$ in $F$ gets identified with a subspace homeomorphic to
$\operatorname{Conf}_{T(v_{1})}(S^{1},\mathfrak{D}_{1}^{\Lambda})\times\cdots\times\operatorname{Conf}_{T(v_{m})}(S^{1},\mathfrak{D}_{m}^{\Lambda})\text{.}$
By Lemma 5.2, we conclude that the connected component of $\varkappa_{T}$ in $F$ is homotopy equivalent to
$\operatorname{Conf}_{T(v_{1})}(S^{1},\mathbb{R}^{2})\times\cdots\times\operatorname{Conf}_{T(v_{m})}(S^{1},\mathbb{R}^{2})\text{.}$
Proof it is a fiber bundle: We claim that $\phi$ makes $E=\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ a fiber bundle over $B=\operatorname{Conf}_{\Lambda}(S^{1},\mathbb{R}^{2})$, with fiber $F$. Let $\kappa\in\operatorname{Conf}_{\Lambda}(S^{1},\mathbb{R}^{2})$, say $\kappa=\{C_{1},\dots,C_{m}\}$. For $\varepsilon>0$, call $\kappa^{\prime}=\{C_{1}^{\prime},\dots,C_{m}^{\prime}\}$ in $B$ an $\varepsilon$-perturbation of $\kappa$ if for each $1\leq i\leq m$ the centers of $C_{i}$ and $C_{i}^{\prime}$ differ by less than $\varepsilon$ and the radii of $C_{i}$ and $C_{i}^{\prime}$ differ by less than $\varepsilon$. Choose $\varepsilon>0$ small enough that for all $i\neq j$ the centers of $C_{i}$ and $C_{j}$ are farther than distance $2\varepsilon$ apart. This ensures that each circle in an $\varepsilon$-perturbation $\kappa^{\prime}$ is unambiguously associated in this way to a unique circle in $\kappa$. Given any such $\kappa^{\prime}=\{C_{1}^{\prime},\dots,C_{m}^{\prime}\}$, with $C_{i}^{\prime}$ “close to” $C_{i}$ for each $i$ in the above sense, let $\alpha_{i}^{\kappa^{\prime}}$ be the affine transformation of $\mathbb{R}^{2}$ that scales and translates taking $C_{i}^{\prime}$ to $\mathfrak{C}_{i}^{\Lambda}$. This is uniquely determined by $\kappa^{\prime}$, since it must scale to change the radius of $C_{i}^{\prime}$ to that of $\mathfrak{C}_{i}^{\Lambda}$, and then translate the center of $C_{i}^{\prime}$ to the center of $\mathfrak{C}_{i}^{\Lambda}$.
Let $U\subseteq B$ be the open neighborhood of $\kappa$ consisting of all $\varepsilon$-perturbations of $\kappa$, so $\phi^{-1}(U)$ consists of all configurations in $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ whose top circles form an $\varepsilon$-perturbation of $\kappa$. Let $\psi\colon\phi^{-1}(U)\to F$ be the map that takes a configuration, say with configuration of top circles $\kappa^{\prime}=\{C_{1}^{\prime},\dots,C_{m}^{\prime}\}$ such that $C_{i}^{\prime}$ is close to $C_{i}$ in the above sense, and for each $i$ applies $\alpha_{i}^{\kappa^{\prime}}$ to $C_{i}^{\prime}$ and all circles nested in $C_{i}^{\prime}$. The output of this procedure is a configuration whose top circles are $\mathfrak{C}_{1}^{\Lambda},\dots,\mathfrak{C}_{m}^{\Lambda}$, i.e., an element of $F$. We can subsequently delete these top circles, if we want to view $F$ in that way. Note that $\psi$ is continuous since the $\alpha_{i}^{\kappa^{\prime}}$ are all continuous, and they vary continuously with $\kappa^{\prime}$.
Let $h=\phi\times\psi\colon\phi^{-1}(U)\to U\times F$, so $h$ is continuous, and we claim it is a homeomorphism. We construct the inverse map directly. An element of $U\times F$ is an ordered pair $(\kappa^{\prime},\widehat{\kappa})$, whose first entry is an $\varepsilon$-perturbation $\kappa^{\prime}=\{C_{1}^{\prime},\dots,C_{m}^{\prime}\}$ of $\kappa$, and whose second entry $\widehat{\kappa}$ is the result of taking a configuration with top circles $\mathfrak{C}_{1}^{\Lambda},\dots,\mathfrak{C}_{m}^{\Lambda}$ and deleting these top circles. Let $g\colon U\times F\to\phi^{-1}(U)$ send $(\kappa^{\prime},\widehat{\kappa})$ to the configuration whose top circles are $C_{1}^{\prime},\dots,C_{m}^{\prime}$ and that otherwise consists of, for each $i$, the result of taking the circles of $\widehat{\kappa}$ nested in $C_{i}$ and applying $(\alpha_{i}^{\kappa^{\prime}})^{-1}$ to them, so that they become nested in $C_{i}^{\prime}$. Note that $g$ is continuous. We have that $h\circ g$ and $g\circ h$ are the identity maps by construction, so they are homeomorphisms. Finally, postcomposing $h$ with the projection $U\times F\to U$ yields $\phi$ by construction, so we are done.
@@
Asphericity: Note that $\operatorname{Conf}_{\Lambda}(S^{1},\mathbb{R}^{2})=\operatorname{Conf}_{m}(D^{2},\mathbb{R}^{2})$ since all the non-root vertices of $\Lambda$ are children of the root. By Lemma 2.6, this is homotopy equivalent to $\operatorname{Conf}_{m}(*,\mathbb{R}^{2})$. In particular it is aspherical, so from the long exact sequence we get that $\pi_{k}(F,\varkappa_{T})\cong\pi_{k}(\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2}),\varkappa_{T})$ for all $k\geq 2$. But the connected component of $F$ containing $\varkappa_{T}$ is homotopy equivalent to $\operatorname{Conf}_{T(v_{1})}(S^{1},\mathbb{R}^{2})\times\cdots\times\operatorname{Conf}_{T(v_{m})}(S^{1},\mathbb{R}^{2})$, which by induction is a product of aspherical spaces, hence is aspherical, so we conclude that $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ is aspherical.
@@
Computation of fundamental group: Since $\operatorname{Conf}_{\Lambda}(S^{1},\mathbb{R}^{2})$ is aspherical, we have an exact sequence
$0\to\pi_{1}(F,\varkappa_{T})\to\pi_{1}(\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2}),\varkappa_{T})\to\pi_{1}(\operatorname{Conf}_{\Lambda}(S^{1},\mathbb{R}^{2}),\varkappa_{\Lambda})\to\pi_{0}(F,\varkappa_{T})\to 0\text{.}$
Here we have used the fact that $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ is connected. We also know that
$\pi_{1}(\operatorname{Conf}_{\Lambda}(S^{1},\mathbb{R}^{2}),\varkappa_{\Lambda})\cong B_{m}\text{,}$
where intuitively an element is represented by the circles in $\varkappa_{\Lambda}$ braiding around in the plane and ending up back where they were set-wise (if not circle-wise). Such an element is in the image of $\pi_{1}(\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2}),\varkappa_{T})$ if and only if each top circle ends up at a top circle whose corresponding child of the root has the same type. Thus, viewing $\pi_{1}(\operatorname{Conf}_{\Lambda}(S^{1},\mathbb{R}^{2}),\varkappa_{\Lambda})$ as $B_{m}$, the image of the map $\pi_{1}(\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2}),\varkappa_{T})\to\pi_{1}(\operatorname{Conf}_{\Lambda}(S^{1},\mathbb{R}^{2}),\varkappa_{\Lambda})$ is $B_{m}^{\Pi}$. Here $\Pi$ is the partition of $\{1,\dots,m\}$ where $i$ and $j$ share a block if and only if $v_{i}$ and $v_{j}$ have the same type.

### v2

@@ 5. Fundamental groups of configuration spaces of circles
In this section we prove our main results. First we focus on the unlabeled case.
@@ 5.1. The unlabeled case
Let $T$ be a tree with $n$ non-root vertices. Then $\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$ is aspherical, and its fundamental group is isomorphic to $\operatorname{bAut}(T)$, hence it is a $K(\operatorname{bAut}(T),1)$.
@@
We should mention that sometimes the definition of “K(G,1)K(G,1)” includes a requirement that the space be homotopy equivalent to a CW-complex. It is clear that our spaces have this property. Indeed, $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ is semialgebraic by Proposition 2.4 and hence triangulable by [Hir75], and $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$ is the orbit space of this under a linear action of the finite group $S_{n}$, hence is triangulable, for example by [PS85]. The same therefore holds for each connected component $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ and $\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$.
@@
Let $D$ be an open disk in the plane. Then $\operatorname{UConf}_{n}(S^{1},D)$ and $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$ are homotopy equivalent. For $T$ a tree with $n$ non-root vertices, setting $\operatorname{UConf}_{T}(S^{1},D)\coloneqq\operatorname{UConf}_{n}(S^{1},D)\cap\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$, we also have a homotopy equivalence from $\operatorname{UConf}_{T}(S^{1},D)$ to $\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$. Analogous statements hold in the labeled case.
@@
Without loss of generality $D$ is centered at the origin. Denote the radius of $D$ by $R$. We will prove that both $\operatorname{UConf}_{n}(S^{1},D)$ and $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$ deformation retract onto a certain subspace. For $\kappa=\{C_{1},\dots,C_{n}\}$ in $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$, let $\delta(\kappa)$ be the maximum distance in $\mathbb{R}^{2}$ from some point in some $C_{i}$ to the origin, so $\delta(\kappa)>0$ and $\delta(\kappa)$ varies continuously with $\kappa$. Let $X$ be the subspace of configurations $\kappa$ with $\delta(\kappa)=R/2$. At this point we have $X\subseteq\operatorname{UConf}_{n}(S^{1},D)\subseteq\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$, and we want to deformation retract both spaces onto $X$. For $t\in[0,1]$ let $f_{t}\colon\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})\to\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$ send $\kappa=\{C_{1},\dots,C_{n}\}$ to the configuration
$\left\{\left((1-t)+t\frac{R/2}{\delta(\kappa)}\right)C_{1},\dots,\left((1-t)+t\frac{R/2}{\delta(\kappa)}\right)C_{n}\right\}\text{.}$
The family of maps $f_{t}$ is a straight line homotopy, such that $f_{0}$ is the identity, each $f_{t}$ is the identity on $X$, and the image of $f_{1}$ lies in $X$, hence this is a strong deformation retraction from $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$ onto $X$. The same family of maps also gives a strong deformation retraction from $\operatorname{UConf}_{n}(S^{1},D)$ onto $X$, and so we conclude that $\operatorname{UConf}_{n}(S^{1},D)$ and $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$ are homotopy equivalent. The same argument works in the presence of a tree $T$ and/or in the labeled case, since scaling does not change the property of one circle being nested in another, nor any chosen labeling of the circles. ∎
A key tool in the proof of Theorem A will be fiber bundles. For completeness let us recall the definition now. Background can be found for example in [Hat02, Pages 375–377]. Let $\phi\colon E\to B$ be a continuous surjection of topological spaces and $F$ a topological space. We call $E$ a fiber bundle over $B$ (with respect to $\phi$, and with fiber $F$) if for all $b\in B$ there exists an open neighborhood $U$ of $b$ and a homeomorphism $h\colon\phi^{-1}(U)\to U\times F$ such that the projection $U\times F\to U$ composed with $h$ equals the restriction of $\phi$ to $\phi^{-1}(U)$. Equivalently, there exists a map $\psi\colon\phi^{-1}(U)\to F$ such that $\phi\times\psi$ restricts to a homeomorphism $\phi^{-1}(U)\to U\times F$. Intuitively, $E$ being a fiber bundle over $B$ with fiber $F$ means that $E$ locally resembles $B\times F$. Note that for each $b\in B$ (and any choice of $U$), the homeomorphism $h$ must take $\phi^{-1}(b)$ homeomorphically to $F_{b}\coloneqq\{b\}\times F$. In particular, all the $\phi^{-1}(b)$ are homeomorphic to each other, and to $F$. If we have a fixed basepoint $b_{0}\in B$, we will often identify $F$ with $F_{b_{0}}$, and so refer to $F$ as a subspace of $E$.
@@
Now we can begin proving our main result, Theorem A, that $\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$ is aspherical and its fundamental group is isomorphic to $\operatorname{bAut}(T)$. As a remark, it is somewhat more traditional to use fiber bundle arguments when dealing with labeled configurations, for example in the original Fadell–Neuwirth paper [FN62a]. In our situation however, it will turn out to be much easier to deduce the labeled case from the unlabeled case than vice versa, and so we will start with the unlabeled case.
If $n=0$ then the theorem follows from the trivial fact that a single point is contractible, so from this point on we assume $n\geq 1$.
Construction 5.3 (Construction of the map $\phi$).
Given an unlabeled configuration $\{C_{1},\dots,C_{n}\}$ in $\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$, call $C_{i}$ a top circle if it is not nested in any $C_{j}$, i.e, if $\mathbb{R}^{2}\vdash C_{i}$. The number of top circles in any configuration in $\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$ equals the number of children of the root $\varnothing$ of $T$, call this $m$. Since $n\geq 1$, we know $m\geq 1$. Let $\Lambda$ be the induced subtree of $T$ consisting of just the root and its children. If $\{v_{1},\dots,v_{m}\}$ is the set of children of the root, then $T$ is the union of $\Lambda$ with all the $T(v_{i})$ for $1\leq i\leq m$.
Define a surjective map
$\phi\colon\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})\to\operatorname{UConf}_{\Lambda}(S^{1},\mathbb{R}^{2})$
by sending a configuration to the configuration consisting only of its top circles. Set $b_{0}=\underline{\kappa}_{\Lambda}$ and $x_{0}=\underline{\kappa}_{T}$, so $\phi(x_{0})=b_{0}$, and let $F$ consist of all configurations in $\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$ whose top circles are precisely $\underline{C}_{1}^{\Lambda},\dots,\underline{C}_{m}^{\Lambda}$, so $F$ is the fiber of $b_{0}$. Identify $F$ with its image in $\operatorname{UConf}_{n-m}(S^{1},\mathbb{R}^{2})$ obtained by deleting these top circles. Writing $\underline{D}_{i}^{\Lambda}$ for the open disk whose boundary is $\underline{C}_{i}^{\Lambda}$, under this identification the connected component of $\underline{\kappa}_{T}$ in $F$ is identified with a subspace homeomorphic to
$\operatorname{UConf}_{T(v_{1})}(S^{1},\underline{D}_{1}^{\Lambda})\times\cdots\times\operatorname{UConf}_{T(v_{m})}(S^{1},\underline{D}_{m}^{\Lambda})\text{.}$
Proposition 5.4.
The map $\phi$ makes $E=\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$ a fiber bundle over $B=\operatorname{UConf}_{\Lambda}(S^{1},\mathbb{R}^{2})$, with fiber $F$.
Proof.
Fix $\kappa\in\operatorname{UConf}_{\Lambda}(S^{1},\mathbb{R}^{2})$, say $\kappa=\{C_{1},\dots,C_{m}\}$. Choose a path in $\operatorname{UConf}_{\Lambda}(S^{1},\mathbb{R}^{2})$ from $\kappa$ to $\underline{\kappa}_{\Lambda}$, and note that this path induces a bijection from $\{C_{1},\dots,C_{m}\}$ to $\{\underline{C}_{1}^{\Lambda},\dots,\underline{C}_{m}^{\Lambda}\}$. Choose the subscript notation so that this bijection is $C_{i}\mapsto\underline{C}_{i}^{\Lambda}$. Choose $\varepsilon>0$ such that for all $i\neq j$ the centers of $C_{i}$ and $C_{j}$ are farther than distance $2\varepsilon$ apart. Call $\kappa^{\prime}\in\operatorname{UConf}_{\Lambda}(S^{1},\mathbb{R}^{2})$ an $\varepsilon$-perturbation of $\kappa$ if for each $C^{\prime}\in\kappa^{\prime}$ there exists $C_{i}\in\kappa$ such that the centers of $C_{i}$ and $C^{\prime}$ differ by less than $\varepsilon$ and the radii of $C_{i}$ and $C^{\prime}$ differ by less than $\varepsilon$. By our choice of $\varepsilon$, each $C^{\prime}$ in $\kappa^{\prime}$ satisfies this condition for a unique $i$, and the function $\kappa^{\prime}\to\kappa$ sending $C^{\prime}$ to this $C_{i}$ is a bijection. Denote by $C_{i}^{\prime}$ the element of $\kappa^{\prime}$ associated to $C_{i}$ under this bijection, and from now on when we write an $\varepsilon$-perturbation of $\kappa$ as $\kappa^{\prime}=\{C_{1}^{\prime},\dots,C_{m}^{\prime}\}$, the above bijection is understood to be $C_{i}^{\prime}\mapsto C_{i}$. Note that this also yields a fixed bijection $\kappa^{\prime}\to\underline{\kappa}_{\Lambda}$ given by $C_{i}^{\prime}\mapsto\underline{C}_{i}^{\Lambda}$. Given an $\varepsilon$-perturbation $\kappa^{\prime}$ of $\kappa$, let $\alpha_{i}^{\kappa^{\prime}}$ be the affine transformation of $\mathbb{R}^{2}$ that scales and translates taking $C_{i}^{\prime}$ to $\underline{C}_{i}^{\Lambda}$. This is uniquely determined by $\kappa^{\prime}$, since it must scale to change the radius of $C_{i}^{\prime}$ to that of $\underline{C}_{i}^{\Lambda}$ and then translate the center of $C_{i}^{\prime}$ to the center of $\underline{C}_{i}^{\Lambda}$, and so the $\alpha_{i}^{\kappa^{\prime}}$ vary continuously with $\kappa^{\prime}$ in the space of affine transformations of $\mathbb{R}^{2}$.
Let $U\subseteq B$ be the open neighborhood of $\kappa$ consisting of all $\varepsilon$-perturbations of $\kappa$, so $\phi^{-1}(U)$ consists of all configurations in $\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$ whose top circles form an $\varepsilon$-perturbation of $\kappa$. Let $\psi\colon\phi^{-1}(U)\to F$ be the map that takes a configuration, say with configuration of top circles the $\varepsilon$-perturbation $\kappa^{\prime}=\{C_{1}^{\prime},\dots,C_{m}^{\prime}\}$, and for each $1\leq i\leq m$ applies $\alpha_{i}^{\kappa^{\prime}}$ to $C_{i}^{\prime}$ and all circles in the configuration that are nested in $C_{i}^{\prime}$. The output of this procedure is a configuration whose top circles are $\underline{C}_{1}^{\Lambda},\dots,\underline{C}_{m}^{\Lambda}$, i.e., an element of $F$. Note that $\psi$ is continuous, since the $\alpha_{i}^{\kappa^{\prime}}$ are all individually continuous, and they vary continuously with $\kappa^{\prime}$.
[…]
@@
Our goal is to prove that $\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$ is aspherical and its fundamental group is isomorphic to $\operatorname{bAut}(T)$. We retain the notation from the proof of Proposition 5.4. Note that $\operatorname{UConf}_{\Lambda}(S^{1},\mathbb{R}^{2})\cong\operatorname{UConf}_{m}(D^{2},\mathbb{R}^{2})$ by Lemma 4.7, so by Lemma 2.6, it is homotopy equivalent to $\operatorname{UConf}_{m}(*,\mathbb{R}^{2})$. In particular it is aspherical, so from the long exact sequence we get that $\pi_{k}(F,\underline{\kappa}_{T})\cong\pi_{k}(\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2}),\underline{\kappa}_{T})$ for all $k\geq 2$. The connected component of $F$ containing $\underline{\kappa}_{T}$ is homotopy equivalent to $\operatorname{UConf}_{T(v_{1})}(S^{1},\mathbb{R}^{2})\times\cdots\times\operatorname{UConf}_{T(v_{m})}(S^{1},\mathbb{R}^{2})$ by Lemma 5.2, which by induction is a product of aspherical spaces, hence is aspherical, so we conclude that $\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$ is aspherical.
@@
Now we compute the fundamental group. Since $\operatorname{UConf}_{\Lambda}(S^{1},\mathbb{R}^{2})$ is aspherical, we have an exact sequence
$0\to\pi_{1}(F,\underline{\kappa}_{T})\to\pi_{1}(\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2}),\underline{\kappa}_{T})\to\pi_{1}(\operatorname{UConf}_{\Lambda}(S^{1},\mathbb{R}^{2}),\underline{\kappa}_{\Lambda})\to\pi_{0}(F,\underline{\kappa}_{T})\to 0\text{.}$
Here we have used the fact that $\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$ is connected. As we said previously, $\operatorname{UConf}_{\Lambda}(S^{1},\mathbb{R}^{2})\simeq\operatorname{UConf}_{m}(*,\mathbb{R}^{2})$, so $\pi_{1}(\operatorname{UConf}_{\Lambda}(S^{1},\mathbb{R}^{2}),\underline{\kappa}_{\Lambda})\cong B_{m}$. An element of this group is represented by a path $\gamma(t)$ ($0\leq t\leq 1$) in $\operatorname{UConf}_{\Lambda}(S^{1},\mathbb{R}^{2})$ from $\underline{\kappa}_{\Lambda}$ to itself. The standard map from $B_{m}$ to the symmetric group $S_{m}$ assigns to $\gamma(t)$ a bijection $\sigma_{\gamma}$ from $\underline{\kappa}_{\Lambda}$ to itself. For each $1\leq i\leq m$ let $C_{i}(t)$ be the path of circles from $\underline{C}_{i}^{\Lambda}$ to $\underline{C}_{\sigma_{\gamma}(i)}^{\Lambda}$ induced by $\gamma$, so $\gamma(t)=\{C_{1}(t),\dots,C_{m}(t)\}$ for all $0\leq t\leq 1$. For each $1\leq i\leq m$ let $\underline{\kappa}_{T(v_{i})}^{C_{i}(t)}$ be the image of $\underline{\kappa}_{T(v_{i})}$ under the map that scales and translates the unit disk to the disk bounded by $C_{i}(t)$. Define $\widetilde{\gamma}(t)=\gamma(t)\cup\bigcup_{i=1}^{m}\underline{\kappa}_{T(v_{i})}^{C_{i}(t)}$, so by Construction 3.3 $\widetilde{\gamma}(t)$ is a path in $\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$ starting at $\widetilde{\gamma}(0)=\underline{\kappa}_{T}$. The element of $B_{m}$ represented by $\gamma(t)$ lies in the image of $\pi_{1}(\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2}),\underline{\kappa}_{T})\to\pi_{1}(\operatorname{UConf}_{\Lambda}(S^{1},\mathbb{R}^{2}),\underline{\kappa}_{\Lambda})$ if and only if $\widetilde{\gamma}(1)=\underline{\kappa}_{T}$. This happens if and only if $v_{i}$ and $v_{\sigma_{\gamma}(i)}$ have the same type for each $1\leq i\leq m$. We conclude this image is $B_{m}^{\Pi}$, where $\Pi$ is the partition of $\{1,\dots,m\}$ in which $i$ and $j$ share a block if and only if $v_{i}$ and $v_{j}$ have the same type.